# The Harmonic Exponential Filter for Nonparametric Estimation on Motion Groups

### v1

@@ III-C Bayesian Filtering
Following the notation presented by Thrun et al. [1, Chapter 2], equations (5) and (6) allow us to incorporate the harmonic exponential distribution into the Bayesian filtering framework:
@@
Computing the predicted belief is commonly referred to as the “prediction step” and requires one to define the motion model. The motion model is the likelihood of the current pose $x_{t}\in SE(N)$ given the previous pose $x_{t-1}\in SE(N)$ and control input $u_{t}$, that is, $p(x_{t}|u_{t},x_{t-1})$. If we assume that the motion is fully defined by the control input, the motion model becomes the likelihood of the relative motion from $x_{t-1}$ to $x_{t}$, or $x^{-1}_{t-1}\circ x_{t}$ matching the relative motion induced by $u_{t}$ or $p_{u_{t}}(x^{-1}_{t-1}\circ x_{t})$. Our prediction step becomes
@@
Input: Previous posterior belief $bel_{t-1}$, relative motion likelihood $p_{u_{t}}$, current measurement likelihood $p_{z_{t}|x_{t}}$ and IUR $T$.
@@
3:// Fourier coefficients
4:$\eta_{z_{t}|x_{t}}\leftarrow\mathcal{F}[\ln(p_{z_{t}|x_{t}})]$triangleright Eq. (4)
5:$\eta_{b\overline{el}_{t}}\leftarrow\mathcal{F}[\ln(\overline{bel_{t}})]$triangleright Eq. (4)
@@
7:$bel(x_{t})=\gamma\exp((\eta_{z_{t}|x_{t}}+\eta_{b\overline{el}_{t}})\cdot T(x_{t}))$triangleright Eq. 9
8:return $bel_{t}$
The update step of a Bayes’ filter is the product of two probability distributions, $p(z_{t}|x_{t})\overline{bel}(x_{t})$. If we perform Fourier analysis in the domain $x_{t}$ of the measurement log-likelihood function we can get a set of Fourier coefficients $\eta_{z_{t}|x_{t}}\triangleq\mathcal{F}[\text{ln}(p(z_{t}|x_{t})]$, and the measurement likelihood function becomes $p(z_{t}|x_{t})=\frac{1}{Z_{\eta}}\exp(\eta_{z_{t}|x_{t}}\cdot T(x_{t}))$. With that and using (5), the update step simplifies to
$bel(x_{t})=\gamma\exp((\eta_{z_{t}|x_{t}}+\eta_{b\overline{el}_{t}})\cdot T(x_{t})),$ (9)
where the normalizing constant $\gamma$ is computed from the zeroth Fourier coefficient as shown in (4).
From this point onward, we use the notation $bel_{t}:SE(N)\to\mathbb{R}^{\geq 0}$ to denote the probability density function at time $t$ associated with the posterior distribution. The predicted belief $\overline{bel}_{t}$, motion model $p_{u_{t}}$, and measurement likelihood $p_{z_{t}|x_{t}}$ are denoted similarly. Note that the previous definitions are functions and not point estimates. Therefore, we can compute the Fourier coefficients of our density functions as $\eta_{z_{t}|x_{t}}\triangleq\mathcal{F}[\ln(p_{z_{t}|x_{t}})]$ and $\eta_{b\overline{el}_{t}}\triangleq\mathcal{F}[\ln(\overline{bel_{t}})]$. A detailed solution to the Bayes’ filter using the harmonic exponential distributions is presented in Algorithm 1.
@@ III-D Analysis of the Harmonic Exponential Filter
We would now like to point out several advantages inherent to the formulation of the HEF. One is the efficiency compared to a histogram-based approach. The HEF requires a transition between two different Fourier spaces to calculate the convolution result, which inherently adds complexity. Nevertheless, it can still be more efficient than a standard convolution implementation in 3D. If we consider a discretized space of $n\times n\times n$ in 3D, a convolution has complexity $O(n^{6})$, while a 3D FFT is of $O(n^{3}\log^{3}{n})$, and the tensor multiplication needed for (6) would be $O(n^{4})$.
@@
TABLE I: A runtime comparison between a histogram-based approach and the HEF for computing the convolution of two functions. Hist (PP) is a direct convolution written in pure python. Hist (Scipy) uses the scipy.signal.convolve() function, and HEF (ours) is the approach we presented in (6) implemented in python using the Numpy FFT library.
@@
Another advantage of representing the density function with Fourier coefficients instead of the probability bins used in histogram filters is the fidelity with which they can represent signals. When computing the Fourier coefficients for the harmonic exponential distribution, the FFT operations behave as a low-pass filter on the signals, leading to information loss in the distribution. However, binning a probability distribution also causes a loss of information. To compare the two, we computed the Kullback–Leibler (KL) divergence $D_{KL}(P||Q)$ between the two approaches and a mixture of two von Mises distributions on $S^{1}$. This analysis is shown in Fig. 3. The KL divergence is a measure of the distance between two distributions, computed as
$D_{KL}(P||Q)=\int_{\mathcal{X}}p(x)\log\left(\frac{p(x)}{q(x)}\right)dx.$
@@
Figure 4: Ability of the HEF and other classic filtering approaches to represent the banana distribution. (a) shows the prior and motion model, which are modelled with a rectangular and Gaussian distribution, respectively. Note that the EKF also uses a Gaussian prior. (b) presents the posterior distribution of the four different approaches after five prediction steps. Only the HEF and particle filter are able to capture the banana distribution.
where $I_{0}(\cdot)$ is the Bessel function and $\kappa$ plays a role similar to the inverse of the covariance in a Gaussian distribution. We can also define a mixture of two von Mises distributions, that is, $p(g)=\frac{1}{2}p(g|\mu_{1},\kappa_{1})+\frac{1}{2}p(g|\mu_{2},\kappa_{2})$. Our analysis in Fig. 3 shows that harmonic exponential distributions have a smaller $D_{KL}$ across a range of situations. When both approaches use few parameters, the harmonic exponential distribution does a better job of representing the underlying density. Increasing the number of parameters improves the fidelity of both approaches; however, the harmonic exponential distribution maintains its advantage in terms of $D_{KL}$. We can also use the $\kappa$ parameter to show the “low-pass filter” effect of taking the FFT. As $\kappa$ increases, the distribution includes more high-frequency components, and $D_{KL}$ increases, but as Fig. 3 shows, the histogram approach performs even worse.
Another benefit of the HEF is its expressivity in capturing diverse distributions, such as the banana shape distribution first presented by Thurn et al. [1, Chapter 5]. This distribution appears when the uncertainty in the heading of an agent increases due to the accumulation of error during multiple prediction steps. Unimodal approaches such as the EKF fail to capture this distribution as it can only model the uncertainty using error ellipses, which is inconsistent with the density of the banana distribution, as shown in Fig. 4.
We compare the expressivity of our method against three other well-established filtering approaches by taking five prediction steps in a horizontal straight path to the right. Interestingly, only the HEF and particle filter are able to capture the banana distribution. Note that all filters start from a rectangular prior, as shown in Fig. 4(a), except for the EKF due to its Gaussian assumptions. Although histogram filters are highly expressive, they fail to capture the banana distribution, as 3D convolutions fail to model the effect of rotations in $SE(2)$.
@@ IV Evaluation
To evaluate our method, we use one simulated and two real-world datasets where the robot state is represented in $SE(2)$. Our assessment focuses on two main metrics. The first metric is the absolute trajectory error (ATE), calculated using two different estimators of the robot position: the mode and mean of the posterior belief. The ATE, measured in meters, is computed along the entire trajectory. However, as this metric does not fully capture the non-Gaussian nature of multimodal estimators, we also calculate the negative log-likelihood (NLL) of the ground truth position under the current belief. We use the equation $\text{NLL}=-\frac{1}{T}\sum_{t}\log bel(x_{t}^{GT})$ where $T$ is the total number of time steps and $bel(x_{t}^{GT})$ is the posterior belief of the true state position, $x_{t}^{GT}$, computed at time $t$. The result of the HEF is the posterior belief over the state space which can be used in a state-belief planner [28]. NLL helps determine how useful our approach would be when used with these systems.
To compute the FFTs, we define a grid of size $50\times 50\times 32$ in $SE(2)$. This grid consists of $50\times 50$ equally spaced Cartesian samples over the interval $(x,y)\in[-0.5,0.5]^{2}$ and 32 equally spaced angular samples $\theta\in[0,2\pi)$. We transform the map of all experiments to the size of this grid. A standard procedure to compute the mean for a harmonic exponential distribution is to take the gradient of the log partition function $\nabla_{\eta}\log Z_{\eta}=\mathbb{E}\left[T(x)\right]\triangleq\mu\quad$ [29, Section 3.4.3]. However, an analytical expression for $\nabla_{\eta}\log Z_{\eta}$ is difficult to compute since it involves solving an integral. We approximate the mean relying on the same samples used to compute the FFTs, that is, $\mu\approx\sum_{n}bel(x_{n})T(x_{n})$. The mode of the HEF is also approximated using the FFT samples: $x^{*}=\underset{x\in{x_{1},x_{2},\dots,x_{N}}}{\operatorname*{argmax}}\mathcal{F}^{-1}[\eta_{bel}](x)$.
In these assessments, our proposed HEF presented in Sec. III is compared against the following filtering approaches: an EKF implementation that assumes unimodal Gaussian state and error, a histogram filter (HistF), and a particle filter (PF). For the sake of comparison, HEF and HistF use $80.000$ parameters, that is, a $50\times 50\times 32$ grid and PF uses $80.000$ particles. To determine the NLL of the PF we approximate the density function using a histogram whose resolution is the same one used for the HEF and HistF. The mode of the PF is estimated by taking the heaviest weighted particle before resampling. For all our experiments we use as a motion model the one of a differential drive robot corrupted with Gaussian noise. To estimate the motion and measurement noise of each filter, we sweep over different noise values and took those that minimize the NLL.
@@ IV-A Simulation
The first evaluation is performed in a simulation environment with range-only measurements. This environment includes a series of landmarks arranged in a line. The robot travels in a counterclockwise circle around the landmarks. The landmarks alternate in providing a range measurement at each time step, where we assume known correspondence. We chose to simulate range measurements due to their highly non-Gaussian measurement likelihood, which produces a ring around each landmark, as can be seen in Fig. 5. We present the results of this experiment in Table II, where metrics are computed over 10 different random seeds. Overall, our method outperforms the baselines, obtaining the lowest ATE and NLL in this multimodal localization experiment.
@@
ATE w/ mode (m) ATE w/ mean (m) NLL (-)
EKF 0.093 $\pm$ 0.04 0.093 $\pm$ 0.04 228.23 $\pm$ 27.78
HistF 0.114 $\pm$ 0.03 0.102 $\pm$ 0.02 13.407 $\bm{\pm}$ 0.49
PF 0.138 $\pm$ 0.02 0.101 $\pm$ 0.02 4.040 $\pm$ 0.92
HEF (ours) 0.058 $\pm$ 0.02 0.056 $\bm{\pm}$ 0.02 0.906 $\pm$ 0.90
TABLE II: Quantitative results for the range-only simulator. Reported metrics include the absolute trajectory error (ATE) computed using the mode and the mean of the posterior belief, and the negative log-likelihood (NLL). Lower values are better. We report the mean and standard deviation of these metrics over 10 different random seeds.
@@ IV-B UWB Range Dataset
Figure 6: Setup for the UWB and doors datasets. (a) shows the layout used in the two UWB runs with a Clearpath Husky and five ultra-wideband (UWB) beacons. (b) presents a top-down view of the laboratory used in the doors dataset.
@@
((a)) Five beacons.
((b)) Three beacons.
((c)) One beacon.
Figure 7: Top-down view showing the configuration of the beacons used in the UWB experiments. By reducing the number of beacons moving from (a) to (c), the measurement likelihood transitions from Gaussian to highly non-Gaussian.
The motion of the agent was determined by integrating the inertial measurement unit (IMU) and the wheel encoders of the Husky platform. The motion noise was estimated to have a standard deviation of 3.08 cm and 0.57 degrees per step. Finally, the ground truth was provided by a NaturalPoint OptiTrack motion capture system. We repeat each run with a different number of beacons to demonstrate the robustness of our proposed HEF to handle unimodal and multimodal distributions. The layout of the beacons used in the experiments is shown in Fig. 7.
The results of this evaluation are presented in Table III. In all runs and with different beacon configurations, the HEF is comparable to or better than the baseline filters even in the presence of highly multimodal measurement likelihoods. Although the EKF obtained the best ATE among all filters in the first run with one beacon, its NLL is the highest among all filters, indicating that throughout the experiment the filter is highly uncertain about the ground truth position. As all the results of the second run closely resembled those of the first run, they were not included in the table.

### v2

@@ III-C Bayesian Filtering
Following the notation of Thrun et al. [1, Chapter 2], equations (5) and (6) allow us to incorporate the harmonic exponential distribution into a Bayesian filter:
@@
The prediction step involves defining the motion model, which represents the likelihood of the current pose $x_{t}\in SE(N)$ given the previous pose $x_{t-1}\in SE(N)$ and control input $u_{t}$, i.e., $p(x_{t}|u_{t},x_{t-1})$. Defining the motion model as $x_{t}=x_{t-1}\circ u_{t}$, where $u_{t}$ is a random variable with a known density, allows us to compute the relative motion as $x^{-1}_{t-1}\circ x_{t}=u_{t}$. Therefore, the relative motion likelihood is $p_{u_{t}}(x^{-1}_{t-1}\circ x_{t})$, giving us the prediction step
@@
The update step of a Bayes’ filter is the product of two probability distributions, $p(z_{t}|x_{t})\overline{bel}(x_{t})$. If we perform Fourier analysis in the domain $x_{t}$ of the measurement log-likelihood function we can get a set of Fourier coefficients $\eta_{z_{t}|x_{t}}\triangleq\mathcal{F}[\text{ln}(p(z_{t}|x_{t})]$, and the measurement likelihood function becomes $p(z_{t}|x_{t})=1/Z_{\eta}\exp(\eta_{z_{t}|x_{t}}\cdot T(x_{t}))$. With that and using (5), the update step simplifies to
$bel(x_{t})=\gamma\exp\left((\eta_{z_{t}|x_{t}}+\eta_{b\overline{el}_{t}})\cdot T(x_{t})\right),$ (9)
where the normalizing constant $\gamma$ is computed from the zeroth Fourier coefficient as shown in (4b).
We use $bel_{t}:SE(N)\to\mathbb{R}^{\geq 0}$ to represent the probability density function for the posterior at time $t$. The predicted belief $\overline{bel}_{t}$, motion model $p_{u_{t}}$, and measurement likelihood $p_{z_{t}|x_{t}}$ are similarly denoted. Note that these definitions are functions, not point estimates, allowing us to compute Fourier coefficients for our density functions as $\eta_{z_{t}|x_{t}}\triangleq\mathcal{F}[\ln(p_{z_{t}|x_{t}})]$ and $\eta_{\overline{bel}_{t}}\triangleq\mathcal{F}[\ln(\overline{bel_{t}})]$. The harmonic exponential Bayes filter is presented in Algorithm 1.
@@
Input: Previous posterior belief $bel_{t-1}$, motion likelihood $p_{u_{t}}$, measurement likelihood $p_{z_{t}|x_{t}}$ and IUR $T$.
@@
3:// Compute Fourier Coefficients
4:$\eta_{z_{t}|x_{t}}\leftarrow\mathcal{F}[\ln(p_{z_{t}|x_{t}})]$triangleright Eq. (4a)
5:$\eta_{b\overline{el}_{t}}\leftarrow\mathcal{F}[\ln(\overline{bel_{t}})]$triangleright Eq.(4a)
@@
7:$\gamma^{-1}\leftarrow\mathcal{F}\left[\exp\left((\eta_{z_{t}|x_{t}}+\eta_{b\overline{el}_{t}})\cdot T(x_{t})\right)\right]_{00}^{0}$triangleright Eq. (4b)
8:$bel(x_{t})=\gamma\exp\left((\eta_{z_{t}|x_{t}}+\eta_{b\overline{el}_{t}})\cdot T(x_{t})\right)$triangleright Eq. (9)
9:return $bel_{t}$
@@ III-D Analysis of the Harmonic Exponential Filter
We now point out several advantages inherent to the HEF. One is efficiency compared to a histogram-based approach. The HEF requires a transition between two different Fourier spaces to compute a convolution, which inherently adds complexity. Nevertheless, it can still be more efficient than a standard convolution implementation in 3D. If we consider a discretized space of $n_{s}\times n_{s}\times n_{s}$ in 3D, a convolution has complexity $O(n_{s}^{6})$, while a $SE(2)$ Fourier transform using FFT methods is of $O(n_{s}^{3}\log{n_{s}^{3}})$. The previous calculation assumes $m,n\approx n_{L}\approx n_{s}$ and $\lambda\approx n_{\lambda}\approx n_{s}$. Under these assumptions, the tensor product needed for (6) is $O(n_{\lambda}n_{L}^{3})$ and the complexity of (2) is also $O(n_{s}^{3}\log{n_{s}^{3}})$.
@@
TABLE I: Runtime comparison of a histogram-based approach and the HEF for convolving two functions. Hist (PP) is a direct convolution written in pure python, Hist (Scipy) uses scipy.signal.convolve(), and HEF (ours) implements (6) in Python using the NumPy FFT library.
Another advantage of representing density functions with Fourier coefficients over probability bins in histogram filters is the fidelity with which they can represent signals. When computing the Fourier coefficients for the harmonic exponential distribution, FFT operations behave as a low-pass filter on the signals, leading to information loss in the distribution. However, binning a probability distribution also causes a loss of information. To compare the two, we computed the Kullback–Leibler (KL) divergence $D_{KL}(P||Q)=\int_{\mathcal{X}}p(x)\log\left(\frac{p(x)}{q(x)}\right)dx$, between the two approaches and a mixture of two von Mises distributions on the Lie group $S^{1}$. This analysis is shown in Fig. 3, where the KL divergence is used as a measure of distance between two distributions.
@@
where $I_{0}(\cdot)$ is the Bessel function and $\kappa$ plays a role similar to the inverse of the covariance in a Gaussian distribution. We can also define a mixture of two von Mises distributions, that is, $p(g)=\frac{1}{2}p(g|\mu_{1},\kappa_{1})+\frac{1}{2}p(g|\mu_{2},\kappa_{2})$. Our analysis in Fig. 3 shows that harmonic exponential distributions have a smaller $D_{KL}$ across a range of situations. When both approaches use few parameters, the harmonic exponential distribution does a better job of representing the underlying density. Increasing the number of parameters improves the fidelity of both approaches; however, the harmonic exponential distribution maintains its advantage in terms of $D_{KL}$.
The parameter $\kappa$ illustrates the “low-pass filter” effect of the FFT: as $\kappa$ increases, the distribution gains high-frequency components, leading to a rise in $D_{KL}$, though, as Fig. 3 shows, the histogram approach performs worse. If the underlying density were discontinuous, like a sum of Haar wavelets, the histogram approach might outperform the HEF in terms of $D_{KL}$. However, since many robotics problems involve smoothly varying functions, the HEF is better suited to handle continuous densities, as shown in this paper.
@@
Figure 4: Ability of the HEF and other filtering approaches to represent the banana distribution. (a) shows the prior and motion model, which are modeled with a rectangular and Gaussian distribution, respectively. Both EKF and IEKF use a Gaussian prior. (b) presents the posterior distribution of the four approaches after five prediction steps. The HEF, PF, and IEKF are able to capture the banana distribution.
Another benefit of the HEF is its expressivity in capturing diverse distributions, such as the banana shape distribution presented by Thurn et al. [1, Chapter 5]. This distribution appears when the uncertainty in the heading of an agent increases due to the accumulation of error during multiple prediction steps. The standard EKF fails to capture this distribution, as it models uncertainty with error ellipses, which is inconsistent with the density of the banana distribution. Long et al. [2] showed that a Gaussian density on the Lie algebra can effectively represent the banana distribution, as shown in Fig. 4. However, a Gaussian density in the Lie algebra can only model unimodal distributions.
We compare the expressivity of our method against four other well-established filtering approaches by taking five prediction steps in a horizontal straight path to the right. Interestingly, the HEF, particle filter (PF) and IEKF are able to capture the banana distribution. Note that all filters start from a rectangular prior, as shown in Fig. 4(a), except for the EKF and IEKF due to its Gaussian assumptions. Although histogram filters (HistF) are highly expressive, they fail to capture the banana distribution, as 3D convolutions fail to model the effect of rotations in $SE(2)$.
@@ IV Evaluation
To evaluate our method, we use one simulated and two real-world datasets where the robot state is represented in $SE(2)$. Our assessment focuses on two main metrics. The first metric is the absolute trajectory error (ATE), calculated using two different estimators of the robot position: the mode and mean of the posterior belief. The ATE, measured in meters, is computed along the entire trajectory. However, as this metric does not fully capture the non-Gaussian nature of multimodal estimators, we also calculate the negative log-posterior (NLP) of the ground truth position under the posterior belief. We use the equation $\text{NLP}=-\frac{1}{T}\sum_{t}\log bel(x_{t}^{GT})$ where $T$ is the total number of time steps and $bel(x_{t}^{GT})$ is the posterior belief of the true state position, $x_{t}^{GT}$, computed at time $t$. The result of the HEF is the posterior belief over the state space which can be used in a state-belief planner [38]. NLP helps to determine how useful our approach would be when used with these systems.
To compute the FFTs, we define a $50\times 50\times 32$ grid in $SE(2)$, with $50\times 50$ Cartesian samples over $(x,y)\in[-0.5,0.5]^{2}$ and 32 angular samples over $\theta\in[0,2\pi)$. The experiment maps are scaled to fit this grid. All HEF experiments use $\lambda_{\text{max}}=35$, where the inverse transform (2) is approximated using 111 samples, $m=669$, and $n=32$. We approximate the mean belief relying on the same samples used to calculate the FFTs, that is, $\mu\approx\sum_{n}bel(x_{n})x_{n}$. The mode of the HEF is also approximated using the FFT samples: $x^{*}=\underset{x\in{x_{1},x_{2},\dots,x_{N}}}{\operatorname*{argmax}}\mathcal{F}^{-1}[\eta_{bel}](x)$.
In these assessments, our proposed HEF presented in Sec. III is compared against the following filtering approaches: an EKF implementation that assumes unimodal Gaussian state and error, a histogram filter (HistF), and a particle filter (PF). For the sake of comparison, HEF and HistF use $80,000$ parameters, that is, a $50\times 50\times 32$ grid and PF uses $80,000$ particles. To determine the NLP of the PF we approximate the density function using a histogram whose resolution is the same one used for the HEF and HistF. The mode of the PF is estimated by taking the heaviest weighted particle before resampling. For all our experiments we use as a motion model the one of a differential drive robot corrupted with Gaussian noise. To estimate the motion and measurement noise of each filter, we sweep over different noise values and took those that minimize the NLP.
@@ IV-A Simulation
The first evaluation is performed in a simulation environment with range-only measurements. This environment includes landmarks arranged in a line. The robot travels in a counterclockwise circle around the landmarks, which alternate in providing a range measurement at each time step, while assuming known correspondence. We chose to simulate range measurements due to their highly non-Gaussian measurement likelihood, which produces a ring around each landmark. All filters start from an equally-weighted bi-modal prior, as shown in Fig. 5. The EKF uses the weighted mean and covariance of this mixture as its prior. We present the results of this experiment in Table II, where metrics are computed over 10 different random seeds. Overall, our method outperforms the baselines, obtaining the lowest ATE and NLP in this multimodal localization experiment.
@@
ATE w/ mode (m) ATE w/ mean (m) NLP (-)
EKF 0.101 $\pm$ 0.03 0.101 $\pm$ 0.03 271.2 $\pm$ 23.06
HistF 0.151 $\pm$ 0.02 0.132 $\pm$ 0.02 17.04 $\bm{\pm}$ 0.57
PF 0.131 $\pm$ 0.03 0.102 $\pm$ 0.02 5.521 $\pm$ 1.15
HEF (ours) 0.089 $\pm$ 0.03 0.082 $\bm{\pm}$ 0.02 2.242 $\pm$ 1.11
TABLE II: Quantitative results for the range-only simulator. Reported metrics include the absolute trajectory error (ATE) computed using the mode and the mean of the posterior belief, and the negative log-posterior (NLP). Lower values are better. We report the mean and standard deviation of these metrics over 10 different random seeds.
@@ IV-B UWB Range Dataset
Figure 6: Setup for the UWB and doors datasets. (a) shows the layout for the two UWB runs with a Clearpath Husky and five ultra-wideband (UWB) beacons. This image also provides a top-down view of the beacon configuration, showing how the measurement likelihood shifts from Gaussian to highly non-Gaussian as beacons are removed. (b) presents a top-down view of the laboratory used in the doors dataset.
@@
The motion of the agent was determined by integrating the IMU and wheel encoders of the Husky platform. The motion noise was estimated to have a standard deviation of 3.08 cm and 0.57 degrees per step. Finally, the ground truth was provided by a NaturalPoint OptiTrack motion capture system. We repeat each run with a different number of beacons to demonstrate the robustness of our proposed HEF to handle unimodal and multimodal distributions. The layout of the beacons used in the experiments is shown in Fig. 6(a).